### v3

# Multi-fidelity learning for interatomic potentials: Low-level forces and high-level energies are all you need

Abstract
The promise of machine learning interatomic potentials (MLIPs) has led to an abundance of public quantum mechanical (QM) training datasets. The quality of an MLIP is directly limited by the accuracy of the energies and atomic forces in the training dataset. Unfortunately, most of these datasets are computed with relatively low-accuracy QM methods, e.g., density functional theory with a moderate basis set. Due to the increased computational cost of more accurate QM methods, e.g., coupled-cluster theory with a complete basis set extrapolation, most high-accuracy datasets are much smaller and often do not contain atomic forces. The lack of high-accuracy atomic forces is quite troubling, as training with force data greatly improves the stability and quality of the MLIP compared to training to energy alone. Because most datasets are computed with a unique level of theory, traditional single-fidelity learning is not capable of leveraging the vast amounts of published QM data. In this study, we apply multi-fidelity learning to train an MLIP to multiple QM datasets of different levels of accuracy, i.e., levels of fidelity. Specifically, we perform three test cases to demonstrate that multi-fidelity learning with both low-level forces and high-level energies yields an extremely accurate MLIP—far more accurate than a single-fidelity MLIP trained solely to high-level energies and almost as accurate as a single-fidelity MLIP trained directly to high-level energies and forces. Therefore, multi-fidelity learning greatly alleviates the need for generating large and expensive datasets containing high-accuracy atomic forces and allows for more effective training to existing high-accuracy energy-only datasets. Indeed, low-accuracy atomic forces and high-accuracy energies are all that are needed to achieve a high-accuracy MLIP with multi-fidelity learning.
1 Introduction
Atomistic simulations, e.g., molecular dynamics (MD), find applications in many fields of science, including biology, materials science, and chemistry. The predictive capability of MD depends primarily on the accuracy of the forces acting on each atom. Historically, these forces have been calculated using either quantum mechanical (QM) methods, such as density functional theory (DFT), or classical mechanical methods, such as force fields (FFs), also referred to as interatomic potentials. Both QM and FF methods have limitations. Although QM calculations are very accurate, they are also extremely expensive, with a cost that increases rapidly with the number of electrons (typically to at least the third power). In comparison with QM, FFs are significantly less expensive and typically scale linearly with the number of atoms, rather than with the number of electrons. However, FFs are less accurate than QM because they use inflexible functional forms with relatively few parameters that are fit to limited amounts of data. Recently, machine learning interatomic potentials (MLIPs), [1] and especially neural network-based MLIPs, [2] promise a “best-of-both-worlds” scenario — computational costs and scaling similar to classical FFs but accuracy similar to the QM method used to generate the training data. [3]
While several different MLIP architectures exist, [4, 5, 6, 7, 8, 9, 10] the overall quality of an MLIP, both in terms of accuracy and robustness, [11] is primarily predicated on the training dataset. [12] By contrast to classical FFs, which are typically parameterized with limited amounts of experimental data, training datasets for MLIPs consist of atomic positions with their corresponding energies and/or forces computed with the desired QM method. Because neural networks are extremely flexible “universal function approximators”, developing reliable and transferable neural network MLIPs requires enormous amounts of training data. [13] For this reason, many studies focus on developing large and diverse QM datasets. [14, 15, 16, 17, 18, 19, 20, 21]
While the rapid growth in freely available training datasets is extremely beneficial to the greater scientific community, [22] several challenges currently limit the practical usefulness of the plethora of public QM data. One challenge is incompatibility across different datasets. In order to have a smooth MLIP, all energies and/or forces need to be computed with precisely the same QM method (e.g., the same functional, the same basis set, the same energy cut-off). Even calculations performed with the same QM method but with two different codes (or even with two different versions of the same code) can result in non-negligible differences in the energies and/or forces. [23] Because different datasets are not all computed with the same QM level of theory or the same QM software, naively combining public datasets for training an MLIP is extremely risky. Thus, although there are many published datasets, researchers commonly generate a completely new training dataset for the specific system of interest using their preferred QM method and code.
Another challenge limiting the impact of public training data is that most datasets are computed using relatively low-level QM methods, with DFT being the most common method of choice. [22] Although modern functionals can achieve reasonable accuracy given a large enough basis set, most QM datasets are computed with traditional functionals and moderate basis sets. Unfortunately, highly accurate QM methods (e.g., coupled-cluster, CC, [24] with a complete basis set, CBS, extrapolation) [25] are orders of magnitude more computationally expensive. Consequently, the few datasets computed with high-level QM methods are much smaller in terms of both number of calculations and size of systems (i.e., number of atoms). [17]
A related challenge is that several QM codes cannot provide analytical forces for the most accurate QM methods. [26, 27, 28] Although most QM codes can compute numerical forces at any level of theory, these forces are more approximate and more expensive to compute as they require performing numerous energy calculations at slightly perturbed atomic configurations. For these reasons, although some recent efforts have expended significant computational resources to calculate large amounts of CC-level forces, [29] most datasets computed with high-level QM methods do not contain force data. [17, 30, 31, 32, 33] For example, while the ANI-1x dataset contains DFT-level energies and forces for $\approx$4.5M configurations of small organic molecules in vacuo, the ANI-1ccx dataset contains only CC-level energies for a subset of $\approx$460k configurations. [17]
Force data are especially valuable for MLIP development. Because forces are vector quantities associated with each atom, a single QM calculation returns one system energy but 3NN force components, where $N$ is the number of atoms. Forces are also a local property, whereas system energy is a global property. Thus, force data facilitate more direct learning of the relationship between the local atomic environment and the potential energy surface. Furthermore, as forces are the negative gradient of the energy, training to forces helps ensure a “smooth” potential. For these reasons, inclusion of force data [34] (and, by extension, Hessian data) [35, 36] is extremely advantageous for training speed, MLIP accuracy, and MLIP stability in MD simulations. For example, due to both the lack of force data in ANI-1ccx and the fewer number of molecular structures in ANI-1ccx compared to ANI-1x, previous MLIPs trained solely to the ANI-1ccx dataset were more limited in accuracy despite being trained to more expensive CC-level energy data. [37, 38, 39]
[…]
Figure 1: Comparison of different methods capable of training to multiple datasets at different levels of theory. Although meta-learning and multi-fidelity learning can utilize many datasets, for visual simplicity, only three levels of theory are depicted. Low-level data are represented with density functional theory (DFT). Intermediate-level data are represented with Møller-Plesset second-order perturbation theory (MP2). High-level data are represented with coupled-cluster theory ($C$). While some variation exists in the literature as to the implementation of each method, the depictions in this figure are consistent with the transfer learning and $\Delta$-learning work of Smith et al., [37] the meta-learning work of Allen et al., [44] and the multi-fidelity learning work of Jacobson et al. [45]
[…]
2 Methods
All models in this study utilize the “Hierarchically Interacting Particle Neural Network” (HIP-NN) architecture, a state-of-the-art message-passing graph-based neural network MLIP. [53] A detailed overview of HIP-NN is provided in the Appendix Section. The HIP-NN architecture is fundamentally different from the “Accurate NeurAl networK engINe for Molecular Energies” (ANAKIN-ME, or ANI) [5] feed forward neural network architecture used in previous MFL work. [45] Although our primary conclusions regarding the performance of MFL do not depend on the MLIP architecture chosen, this section describes some MFL design decisions specific to HIP-NN.
2.1 Multi-fidelity learning with HIP-NN
Although HIP-NN has never been used previously to predict energy at multiple levels of theory, standard HIP-NN is already capable of predicting multiple outputs. Therefore, no modifications to the HIP-NN code were required to perform MFL. However, training HIP-NN to multiple levels of theory requires several important design decisions. The most important decision is which layers use shared parameters (i.e., parameters trained simultaneously to multiple levels of theory) and which layers use non-shared parameters (i.e., fidelity-specific parameters trained to a single level of theory). In this work, all interaction blocks use shared parameters while all linear output layers use fidelity-specific non-shared parameters (see Figure 2). Therefore, the weights and biases in Equation 4 and Equation 7 are shared parameters while the weights and biases in Equation 8 are the only non-shared parameters (see Appendix).
Figure 2: HIP-NN multi-fidelity architecture with shared parameters for interaction blocks (green) and fidelity-specific non-shared parameters for linear output layers (yellow and orange). $n$ refers to the interaction block and hierarchical order. $t$ refers to the level of theory. $i$ refers to the atom number.
For multi-fidelity learning, the total loss function $(L)$ includes a separate loss term for energy and forces for each level of theory, according to:
$L=\sum_{t}(w_{E,t}L_{E,t}+w_{F,t}L_{F,t}+w_{H,t}L_{H,t})$ (1)
where $t$ denotes the level of theory, $E$ stands for energy, $F$ stands for forces, $L_{E,t}$, $L_{F,t}$, and $L_{H,t}$ are the energy, force, and hierarchicality loss terms for the $t^{\rm th}$ level of theory, respectively, and $w_{E,t}$, $w_{F,t}$, $w_{H,t}$ are the weights for the energy, force, and hierarchicality loss terms for the $t^{\rm th}$ level of theory, respectively. The loss function for single-fidelity learning is obtained when the summation is over one level of theory. In this work, the weights for all contributions are equal, i.e., $w_{E,t}=1$, $w_{F,t}=1$, and $w_{H,t}=1$. However, force weights are zero $(w_{F,t}=0)$ if not training to forces from that level of theory. Several other studies also use an equal energy-to-force weight ratio when training diverse MLIP architectures to the ANI-1x dataset. [54, 55, 56, 57] Although the present work utilizes constant weights throughout training, future work could investigate whether an adaptive loss weighting approach is beneficial for MFL. [58] Specifically, the low-level forces would be assigned greater weights at early epochs while the high-level energies would be assigned greater weights at later epochs, similar to recent work with knowledge distillation. [59]
The $L_{H,t}$ term serves as a regularization term that ensures the MLIP obeys the principle of hierarchicality, namely, that the hierarchical energy contribution to the total energy decreases for increasing hierarchicality, i.e., $E^{n+1}<E^{n}$ (see Figure 2 and Appendix for details).
In this work, the energy and force loss terms are defined as
$L_{E}=RMSE_{E}+MAE_{E}$ (2)
$L_{F}=RMSE_{F}+MAE_{F}$ (3)
where $RMSE$ is the root-mean-square error and $MAE$ is the mean absolute error.
2.2 Datasets
2.2.1 Training
Training is performed using a subset of the original ANI-1x dataset, which consists primarily of near-equilibrium (i.e., non-reactive) structures for stable organic molecules. [17] While the complete ANI-1x dataset contains $\approx$4.5M molecular structures, we utilize a subset of nearly 460k molecular structures that correspond to the same structures as the ANI-1ccx dataset. Both the complete ANI-1x dataset and the ANI-1ccx subset were generated in previous work through active learning. [60] Although the full ANI-1x dataset contains low-level energies and forces for all $\approx$4.5M structures, high-level energies are available only for this reduced subset. Thus, utilizing the reduced dataset of $\approx$460k structures helps avoid the issue of unequal data volumes between different QM methods. Previous studies have trained reliable MLIPs for diverse architectures to both the complete ANI-1x dataset as well as significantly reduced subsets of ANI-1x. [38, 39, 54, 55, 56, 57, 61, 62]
[…]
2.2.2 Testing
Two types of test datasets are used to evaluate the accuracy of our trained models: in-sample and out-of-sample. In-sample test datasets are generated by randomly selecting 10% of the $\approx$460k configurations in the reduced ANI-1x dataset. These $\approx$50k configurations form a held-out dataset that is not included in training or validation. Each in-sample test dataset uses the same fidelity level as the training data.
Out-of-sample test datasets come from the GDB10to13 portion of the COMP6 dataset, [60] which contains approximately 48k conformations of larger molecules (10–13 heavy atoms) from the GDB-11 database [71, 72] and the GDB-13 database. [73] The GDB10to13 test dataset primarily serves as an extensibility test. Forces in the GDB10to13 test dataset are computed using MP2/TZ, while conformer energy differences $(\Delta E_{\rm conf})$ are computed with CCSD(T)*/CBS. To allow for comparison with previous work and to focus on chemically relevant conformers, only conformers within 100 kcal/mol of the minimum-energy conformer are included in the analysis of GDB10to13 forces and $\Delta E_{\rm conf}$.
An additional out-of-sample test dataset contains torsion scan energy differences $(\Delta E_{\rm tors})$ for small drug-like molecules. This test dataset contains the same geometries as Sellers et al. [74] that were previously optimized with MP2 and the split-valence TZ with diffuse and polarization functions (6-311+G**) basis set. [75] However, the energies in our torsion scan test dataset were previously recomputed using the same CCSD(T)*/CBS approach as the ANI-1ccx dataset. We also test our models against the CCSD(T)/CBS energies reported by Sellers et al. (see Supplementary Material), to allow for a fair comparison with other studies. [37, 45, 59, 76] To further test the robustness of our models, we also perform relaxed torsion scans, wherein the positions of all atoms are optimized with a given model subject to the constraint of a fixed dihedral angle (see Supplementary Material).
Performance for each model is measured by the mean absolute error (MAE) and/or root-mean-square error (RMSE) of energies and atomic forces. The RMSE and MAE are reported as an average value with a corresponding 95% confidence interval from an ensemble of eight MLIPs. Specifically, the RMSE and MAE are first computed for each individual ensemble member. Subsequently, the mean and standard deviation are computed from the eight different RMSE and MAE values. Consequently, the average RMSEs and MAEs reported are not based on ensemble-averaged energies and forces. Averaging the energy and forces over an ensemble of MLIPs smooths the potential energy surface and, thereby, significantly reduces the errors. However, by not averaging the predictions prior to computing the errors, it is possible to quantify the ensemble uncertainty in the errors themselves and, thus, determine whether two models are statistically different. All errors reported in this work correspond to test errors, i.e., the errors computed on either the held-out dataset or the GDB10to13 testset. No training or validation errors are reported.
2.3 Models
A comprehensive set of training tests is performed with each level of theory for both single-fidelity (SF) and multi-fidelity (MF) machine learning interatomic potentials (MLIPs). Table 1 gives a summary of all MLIPs trained in this work, as well as the shorthand notation adopted to distinguish between models. The SF-MLIPs allow for a baseline comparison with the MF-MLIPs as well as a benchmark comparison with the published ANI MLIPs trained to the same level of theory. An ensemble consisting of eight MLIPs is trained for each model. See Supplementary Material for model and training details.
Table 1: All MLIPs trained in this work. SF = single-fidelity. MF = multi-fidelity. DZ = DFT/DZ. TZ = DFT/TZ. MP2 = MP2/TZ. $C$ = CCSD(T)*/CBS. E = energies. F = forces. ✓signifies MLIP was trained with data from this level of theory (DFT/DZ, DFT/TZ, MP2/TZ, or CCSD(T)*/CBS) and property (E or F).
DFT/DZ DFT/TZ MP2/TZ CCSD(T)*/CBS
Experiment Name E F E F E E
SF-DZ/E
SF-DZ/EF
SF-TZ/E
SF-TZ/EF
SF-MP2/E
SF-CC/E
MF-DZ/E-TZ/E
MF-DZ/EF-TZ/E
MF-DZ/E-TZ/EF
MF-DZ/EF-TZ/EF
MF-DZ/EF-MP2/E
MF-TZ/EF-MP2/E
MF-DZ/E-CC/E
MF-DZ/EF-CC/E
MF-TZ/E-CC/E
MF-TZ/EF-CC/E
MF-DZ/EF-TZ/EF-CC/E
3 Results
The primary objective of this study is to demonstrate that, even in the absence of high-level forces, MFL can achieve accuracy similar to SF training directly to high-level energies and forces. We perform the following three test cases to achieve this objective.
The first test case of MFL utilizes DFT-level energies and forces computed with two different basis sets, DZ and TZ. The larger TZ basis set provides the high-level data in this test case. Although we do not anticipate significant differences in forces for such a marginal increase in basis set size, this first test case is insightful because the reduced ANI-1x training dataset contains both low-level (DFT/DZ) and high-level (DFT/TZ) atomic forces. Thus, we are able to train an SF-MLIP directly to DFT/TZ energies and forces (SF-TZ/EF). The SF-TZ/EF model serves as a useful benchmark for comparison with the MF-MLIP trained without DFT/TZ forces (MF-DZ/EF-TZ/E).
The second test case of MFL also utilizes DFT/DZ for the low-level data, but instead utilizes MP2/TZ for the high-level data. In this test case, however, the reduced ANI-1x training dataset does not contain forces for MP2/TZ. Therefore, it is not possible to benchmark the MF-DZ/EF-MP2/E MLIP against an SF-MLIP trained directly to MP2/TZ forces (SF-MP2/EF). The benefit of this second test case is that the GDB10to13 test set contains MP2/TZ forces for comparison.
The third test case of MFL utilizes DFT/DZ for low-level data and CCSD(T)*/CBS for high-level data. CC-level forces are not available for either the ANI-1ccx training dataset or the GDB10to13 test dataset. Thus, not only are we not able to train a SF-MLIP directly to CC-level forces for comparison (SF-CC/EF), but we also cannot test the quality of the MLIP predictions against CC-level forces. Instead, we compare the SF-MLIP and MF-MLIP performance for predicting CC-level conformer energy differences for the GDB10to13 dataset.
3.1 Test case one: DFT/DZ and DFT/TZ
We begin by investigating MFL with DFT/DZ and DFT/TZ data. Because forces at both levels of fidelity are available for training, this test case allows for direct validation of MFL performance.
Table 2 reports the energy and force RMSEs for an SF-MLIP trained just to DFT/TZ energies (SF-TZ/E), an SF-MLIP trained to DFT/TZ energies and forces (SF-TZ/EF), an MF-MLIP trained to DFT/DZ energies and DFT/TZ energies (MF-DZ/E-TZ/E), an MF-MLIP trained to DFT/DZ energies and forces and DFT/TZ energies (MF-DZ/EF-TZ/E), and an MF-MLIP trained to DFT/DZ energies and forces and DFT/TZ energies and forces (MF-DZ/EF-TZ/EF). See Supplementary Material for additional validation results relevant to the DFT/DZ and DFT/TZ test case.
Table 2: Comparison of energy and force root-mean-square errors (RMSEs) of DFT/TZ prediction, using single-fidelity (SF) and multi-fidelity (MF) MLIPs trained with DFT/TZ forces and/or energies with/without DFT/DZ forces and/or energies. MFL with low-level (DFT/DZ) forces improved the prediction on the high-level (DFT/TZ) energies and forces compared to SF training to high-level energies only. RMSEs are computed relative to the DFT/TZ energies and forces for the in-sample test set. Error bars represent 95% confidence interval from an ensemble of eight models.
Energy-RMSE (kcal/mol) Force-RMSE (kcal/mol/Å)
SF-TZ/E 2.75 $\pm$ 0.02 13.52 $\pm$ 0.50
SF-TZ/EF 1.60 $\pm$ 0.02 3.52 $\pm$ 0.09
MF-DZ/E-TZ/E 2.78 $\pm$ 0.04 11.92 $\pm$ 0.15
MF-DZ/EF-TZ/E 1.49 $\pm$ 0.04 4.16 $\pm$ 0.15
MF-DZ/EF-TZ/EF 1.34 $\pm$ 0.02 3.42 $\pm$ 0.11
A key observation from Table 2 is that MFL with low-level forces (MF-DZ/EF-TZ/E) clearly outperforms SF training to only high-level energies (SF-TZ/E). Specifically, the energy and force RMSEs for MF-DZ/EF-TZ/E (1.49 kcal/mol and 4.16 kcal/mol/Å) are approximately a factor of two or three times smaller than for SF-TZ/E (2.75 kcal/mol and 13.52 kcal/mol/Å), respectively. These results provide strong evidence that, in the common scenario where high-level forces are not available for training, MFL with low-level forces is far superior to SF training to only high-level energies.
Furthermore, MFL with low-level forces achieves similar accuracy as training directly to high-level forces. Specifically, the MF-DZ/EF-TZ/E force errors (4.16 kcal/mol/Å) are only 18% higher than the SF-TZ/EF force errors (3.52 kcal/mol/Å) and the MF-DZ/EF-TZ/E energy errors (1.49 kcal/mol) are actually 7% lower than the SF-TZ/EF energy errors (1.60 kcal/mol). This performance is quite remarkable considering that MF-DZ/EF-TZ/E was not trained with DFT/TZ forces. Therefore, in the absence of high-level forces, MFL with low-level forces is sufficient for training an MLIP that predicts accurate high-level energies and forces.
By contrast, MFL with only energies at both levels of theory (MF-DZ/E-TZ/E) provides limited improvement compared to SF training without forces (SF-TZ/E). The force RMSE for MF-DZ/E-TZ/E remains too large (approximately 12 kcal/mol/Å) for practical use. The lack of improvement for MFL without low-level forces is consistent with previous work. [45]. Thus, alternative training methods should be investigated if forces are not available at either level of theory.
Although the performance of MFL with low-level forces is impressive, SF training directly to high-level forces is still marginally better for force prediction. If high-level forces are available for training, however, it is also possible to perform MFL with both low-level and high-level forces (MF-DZ/EF-TZ/EF). In fact, the MF-DZ/EF-TZ/EF errors are slightly lower than the SF-TZ/EF errors for both energy (1.34 kcal/mol vs 1.60 kcal/mol) and forces (3.42 kcal/mol/Å vs 3.52 kcal/mol/Å). Therefore, MFL with both low-level and high-level forces outperforms SF learning to each level of theory separately.
3.2 Test case two: DFT/DZ and MP2/TZ
We now investigate MFL with DFT/DZ and MP2/TZ data. Although the training dataset does not contain any MP2/TZ force data, this test case allows for validation against the MP2/TZ forces in the GDB10to13 dataset. In the first test case, the two levels of theory (DFT/DZ and DFT/TZ) are relatively similar and computationally inexpensive. In this test case, however, the MP2/TZ forces are quite different from the DFT/DZ forces (RMSE of 5.9 kcal/mol/Å when comparing DFT/DZ and MP2/TZ forces in GDB10to13). The more expensive MP2/TZ forces are also assumed to be more accurate than DFT/TZ, considering MP2 is a higher level of theory and both methods use similar triple zeta basis sets. Therefore, this test case presents a more interesting and practical challenge for MFL than test case one.
Table 3 reports the energy and force RMSEs for a single-fidelity MLIP trained just to MP2/TZ energies (SF-MP2/E) and a multi-fidelity MLIP trained to DFT/DZ energies and forces and MP2/TZ energies (MF-DZ/EF-MP2/E). See Supplementary Material for additional validation results relevant to the DFT/DZ and MP2/TZ test case.
Table 3: Comparison of energy and force root-mean-square errors (RMSEs) of MP2/TZ prediction, using single-fidelity (SF) and multi-fidelity (MF) MLIPs trained with MP2/TZ energies with/without DFT/DZ forces and energies. MFL with low-level (DFT/DZ) forces improved the prediction on the high-level (MP2/TZ) energies and forces compared to SF training to high-level energies only. No MP2/TZ forces were used during training of MLIPs. Energy RMSEs are computed relative to the MP2/TZ energies for the in-sample test set. Force RMSEs are computed relative to the MP2/TZ forces in the GDB10to13 out-of-sample test set. Error bars represent 95% confidence interval from an ensemble of eight models.
Energy-RMSE (kcal/mol) Force-RMSE (kcal/mol/Å)
SF-MP2/E 2.95 $\pm$ 0.02 8.59 $\pm$ 0.42
MF-DZ/EF-MP2/E 1.72 $\pm$ 0.02 3.42 $\pm$ 0.09
Consistent with the first test case, the results in Table 3 demonstrate that MFL with low-level (DFT/DZ) forces is significantly more accurate than SF training to only high-level (MP2/TZ) energies. Specifically, both the energy and force RMSEs for MF-DZ/EF-MP2/E (1.72 kcal/mol and 3.42 kcal/mol/Å) are approximately a factor of two lower than for SF-MP2/E (2.95 kcal/mol and 8.59 kcal/mol/Å), respectively. Furthermore, the force RMSE for MF-DZ/EF-MP2/E on the GDB10to13 dataset (3.42 kcal/mol/Å) is significantly lower than the RMSE between DFT/DZ forces and MP2/TZ forces (5.9 kcal/mol/Å). Thus, MFL is outperforming the accuracy that could be achieved by simply learning the low-level (DFT/DZ) forces.
3.3 Test case three: DFT/DZ and CCSD(T)*/CBS
Having verified that MFL quantifiably improves the high-level energies and forces for the first two test cases, we now aim at the most challenging and important test case, namely, MFL with DFT/DZ and gold-standard CCSD(T)*/CBS data. Because CC is prohibitively expensive, no CC-level force data are available for direct validation in either the ANI-1ccx training dataset or the GDB10to13 test dataset. However, Supplementary Material provides indirect evidence that MFL with low-level forces improves CC-level forces.
We perform two additional tests using CCSD(T)*/CBS energies for conformers from the GDB10to13 dataset and torsion scans from the Sellers et al. dataset. [74] The conformer energies are reported as energy differences $(\Delta E_{\rm conf})$ between all possible pairs of conformers. The torsion energies are reported as energy differences $(\Delta E_{\rm tors})$ relative to the minimum energy for a given torsion scan. Predicting accurate conformer and torsion energies is essential for determining the most stable structure of a molecule. Conformer geometry searches can be quite expensive when performed with QM methods, whereas classical FFs may not be sufficiently accurate to correctly identify the minimum-energy conformer geometry. [77] Thus, an MLIP that can predict accurate CC-level conformer and torsion energy differences would be extremely valuable for real-life applications, such as drug discovery. [78, 79]
Table 4 reports the energy and $\Delta E_{\rm conf}$ RMSEs for a single-fidelity MLIP trained to CCSD(T)*/CBS energies (SF-CC/E), a multi-fidelity MLIP trained to DFT/DZ energies and CCSD(T)*/CBS energies (MF-DZ/E-CC/E), and a multi-fidelity MLIP trained to DFT/DZ energies and forces and CCSD(T)*/CBS energies (MF-DZ/EF-CC/E). See Supplementary Material for additional validation results relevant to the DFT/DZ and CCSD(T)*/CBS test case.
Table 4: Comparison of energy, conformer energy differences ($\Delta E_{\rm conf}$), and torsion energy differences ($\Delta E_{\rm tors}$) root-mean-square errors (RMSEs) of CCSD(T)*/CBS prediction, using single-fidelity (SF) and multi-fidelity (MF) MLIPs trained with CCSD(T)*/CBS energies with/without DFT/DZ forces and/or energies. MFL with low-level (DFT/DZ) forces improved the prediction on the high-level (CCSD(T)*/CBS) energies compared to SF training to high-level energies only. No CCSD(T)*/CBS forces were used during training of SF or MF MLIPs. Energy RMSEs are computed relative to the CCSD(T)*/CBS energies for the in-sample test set. $\Delta E_{\rm conf}$ and $\Delta E_{\rm tors}$ RMSEs are computed relative to the CCSD(T)*/CBS energy differences for the GDB10to13 conformer and torsion scan out-of-sample test sets, respectively. Error bars represent 95% confidence interval from an ensemble of eight models.
Energy-RMSE (kcal/mol) $\Delta E_{\rm conf}$-RMSE (kcal/mol) $\Delta E_{\rm tors}$-RMSE
SF-CC/E 2.78 $\pm$ 0.05 2.65 $\pm$ 0.04 0.71 $\pm$ 0.03
MF-DZ/E-CC/E 2.76 $\pm$ 0.03 2.60 $\pm$ 0.03 0.67 $\pm$ 0.04
MF-DZ/EF-CC/E 1.52 $\pm$ 0.03 1.60 $\pm$ 0.03 0.40 $\pm$ 0.02
Similar to the results for the previous two test cases, the results in Table 4 demonstrate that MFL with low-level (DFT/DZ) forces is significantly more accurate than SF training to only high-level (CCSD(T)*/CBS) energies. Specifically, the MF-DZ/EF-CC/E errors are considerably lower than the SF-CC/E errors on CCSD(T)*/CBS energies (1.52 kcal/mol vs 2.78 kcal/mol), CCSD(T)*/CBS conformer energy differences (1.60 kcal/mol vs 2.65 kcal/mol), and CCSD(T)*/CBS torsion scan energy differences (0.40 kcal/mol vs 0.71 kcal/mol), respectively. The MF-DZ/EF-CC/E energy, $\Delta E_{\rm conf}$, and $\Delta E_{\rm tors}$ RMSEs (1.52 kcal/mol, 1.60 kcal/mol, and 0.40 kcal/mol) are also substantially lower than the RMSEs for the ANI-1ccx MLIP (2.57 kcal/mol, 2.62 kcal/mol, and 0.77 kcal/mol), respectively. See Supplementary Material for additional comparisons between HIP-NN and ANI models. By contrast, MFL without forces (i.e., only energies for both levels of theory) is essentially as accurate as SF training to just the high-level energies, consistent with previous work. [45] This result demonstrates that including low-level forces is essential for MFL to achieve significant improvement in accuracy.
4 Discussion
These three test cases clearly demonstrate that MFL with low-level forces and high-level energies outperforms SF training with only high-level energies. Moreover, MFL with low-level forces is of similar accuracy as SF training directly to high-level forces, in cases where those are available for comparison. To help quantify the improvement MFL provides in all three test cases, Figure 3 presents the percent decrease in RMSE for energies, forces, and conformer energy differences between each MF-MLIP trained with low-level (DFT/DZ) forces and the corresponding SF-MLIP trained with only high-level (either DFT/TZ, MP2/TZ, or CCSD(T)*/CBS) energies. See Supplementary Material for additional analysis of percent improvement.
[…]
The percent improvement in energy and conformer energy differences with MFL is between 40% and 45% for all three test cases. The percent improvement in forces with MFL is 70% for MF-DZ/EF-TZ/E tested on the ANI-1x DFT/TZ forces and 60% for MF-DZ/EF-MP2/E tested on the GDB10to13 MP2/TZ forces. For comparison, the percent improvement with TL is more than a factor of two smaller than with any of the MFL test cases. Specifically, based on the reported errors for the ANI-1ccx and ANI-1ccx-R MLIPs [37], which were trained with and without TL, respectively, the percent improvement with TL is between 15-20% for energies and conformer energies. For a truly fair comparison between TL and MFL, however, future work could investigate the percent improvement for TL that is achieved when pre-training with the same HIP-NN architecture and the same reduced ANI-1x dataset with both DFT/DZ energies and DFT/DZ forces.
One unexpected result is that MFL leads to a measurable decrease in accuracy on the low-level forces themselves (see Supplementary Material). Specifically, the RMSEs on the DFT/DZ forces for MF-DZ/EF-TZ/E (3.95 kcal/mol/Å), MF-DZ/EF-MP2/E (3.91 kcal/mol/Å), and MF-DZ/EF-CC/E (3.76 kcal/mol/Å) are all markedly higher than the force RMSE for SF-DZ/EF (3.17 kcal/mol/Å). This increase in error on the low-level (DFT/DZ) forces suggests that the current MFL approach requires a compromise in accuracy on the low-level forces to achieve improved accuracy on the high-level forces. Future work could attempt to elucidate whether this is a fundamental limitation of MFL, a limitation of our implementation, or a limitation in the model architecture. Some possible considerations to help reduce the low-level force errors are to use a larger HIP-NN architecture (i.e., more interactions layers, more atom layers, more features), and/or modify the number of shared parameters, and/or include non-linear layers for the fidelity-specific output nodes. However, we emphasize that the primary goal of this work is to improve the high-level energy and force predictions. Indeed, if only accurate prediction of the low-level energy and forces is the desired outcome, SF learning would suffice.
For completeness, we also investigate the potential benefit of applying MFL with forces from more than one level of theory. In this case, the energies and forces for both DFT/DZ and DFT/TZ are included in training along with the CCSD(T)*/CBS energies. Notably, the CC-level errors are actually considerably worse for MFL with three datasets than MFL with just two datasets (see Supplementary Material). With twice as much DFT-level data, the worse performance with three datasets is likely attributed to a data imbalance. An alternative explanation is that the DFT/DZ and DFT/TZ forces are providing contradictory information that MFL simply cannot resolve. Another possibility is that the MLIP has reached its capacity to match the different datasets with the chosen architecture and hyperparameters. However, with the current MFL approach, low-level forces from just a single level of theory are sufficient to achieve the desired accuracy on the high-level forces.
5 Conclusions and Future Work
In the age where datasets with low-level quantum mechanical forces are in abundance but high-level forces are scarce, multi-fidelity learning is a powerful technique to develop machine learning interatomic potentials with gold-standard coupled-cluster-level accuracy. Although previous studies with multi-fidelity learning to only energy data did not report significant improvements compared to single-fidelity learning, the current work shows that the inclusion of low-level forces is all that is needed to achieve high-level accuracy when the two datasets cover the same configuration space.
The results of all three test cases are complementary, and the conclusions are unanimous. First, multi-fidelity learning with low-level forces achieves a factor of two improvement in high-level energies and forces compared to single-fidelity learning with only high-level energies. Second, multi-fidelity learning with low-level forces achieves similar accuracy as single-fidelity learning directly to high-level forces. Third, multi-fidelity learning with both low-level and high-level forces achieves the highest accuracy. Fourth, multi-fidelity learning without forces does not provide significant benefit compared to single-fidelity learning with only energies.
While the present study focused on multi-fidelity learning with just two or three quantum mechanical datasets, multi-fidelity learning is capable of leveraging several datasets with vastly different amounts of data. Future work could investigate alternative multi-fidelity learning training protocols that better account for data volume imbalance between different levels of theory. This study considered multi-fidelity learning when both the low-level and high-level data exist for the same molecular structures. While this approach clearly improves accuracy on the types of systems included in the training set, it is unlikely to improve transferability. Future work could consider how multi-fidelity learning can leverage disparate datasets to share knowledge from different regions of chemical space. Ideally, the molecular structures would cover completely different portions of the potential energy surface, although some overlap may be necessary for joint learning. [51] For example, future work could determine if multi-fidelity learning with low-level forces and high-level energies is beneficial when training to both non-reactive (e.g., ANI-1ccx) and reactive (e.g., Transition1x) [18] datasets or when training to both in vacuo (e.g., ANI-1ccx) and periodic boundary (e.g., high-pressure MD) [80, 81, 82] datasets. Future work could also test the limits of multi-fidelity learning by using low-level forces that are computed with extremely inexpensive but even more approximate methods, such as classical force fields or semi-empirical quantum chemistry. [83] Future work could also consider including in the multi-fidelity dataset so-called “atomic energies.” Although atomic energies are not physical quantities that can be computed with quantum mechanical methods, previous studies have shown the advantage of (pre-)training to atomic energies [84, 85] computed with either a classical force field [86] or a “teacher” machine learning interatomic potential. [59, 87] Furthermore, in light of the recent advances training with Hessian data, [35, 36] future work could also investigate multi-fidelity learning with low-level Hessians.
Data availability
The data that support the findings of this study will be openly available following an embargo at the following URL/DOI: https://github.com/lanl/hippynn/examples/multifidelity_ani1x_training.py.
Acknowledgments
M.M., S.M., A.E.A.A., B.N., K.B., N.L., and R.A.M. acknowledge support from the US Department of Energy, Office of Science, Basic Energy Sciences, Chemical Sciences, Geosciences, and Biosciences Division under Triad National Security, LLC (‘Triad’) contract grant 89233218CNA000001 (FWP: LANLE3F2 and LANLE8AN). M.M. gratefully acknowledges the resources of the Los Alamos National Laboratory (LANL) Computational Science summer student program. The work at LANL was supported by the LANL Laboratory Directed Research and Development (LDRD) Project 20230290ER. Work at LANL was performed in part at the Center for Nonlinear Studies and the Center for Integrated Nanotechnologies, a US Department of Energy Office of Science user facility at LANL. This research used resources provided by the Darwin testbed at LANL which is funded by the Computational Systems and Software Environments subprogram of LANL’s Advanced Simulation and Computing program (NNSA/DOE). This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.
Appendix
Terminology
Although the terms multi-fidelity, multi-level, multi-task, and multi-head are used somewhat interchangeably in the literature, we briefly explain some subtle differences between these methods within the context of MLIP development.
Typical MLIPs predict only a single energy. To ensure conservation of energy, forces are not, or at least “should” not be, predicted directly as a second output. Rather, the MLIP forces are obtained by differentiation of the energy with respect to the atomic positions. Thus, most MLIPs possess a single output node, often referred to as a “head” in the context of neural networks. Although the loss function might include multiple error terms that depend on energies and forces (and sometimes stresses), these error terms do not require separate training “tasks”. Thus, traditional training is neither multi-task nor multi-head learning, although it may still be considered a multi-objective optimization.
Multi-head learning is performed whenever an MLIP directly predicts multiple outputs. These multiple output heads can predict the energy for different QM levels of theory, as in the case of multi-fidelity (or multi-level) learning. However, in addition to predicting energies, these output heads can also predict atomic or molecular properties (e.g., dipole moment, electronegativity). [49, 50, 88] As learning these multiple properties requires separate training “tasks”, this example of multi-head learning is also referred to as multi-task learning. However, as each property is typically computed using a single QM level of theory, this is not an example of multi-fidelity learning. Training to energies and forces at multiple levels of theory also requires separate training “tasks.” Therefore, although multi-head learning and multi-task learning are not necessarily the same as multi-fidelity learning, the multi-fidelity learning approach used in this study can also be considered multi-level, multi-head, and multi-task learning.
Not all multi-fidelity learning approaches utilize multi-head learning. For example, an alternative multi-fidelity approach is to provide the desired level of theory directly to the neural network as an input, such that a single output head predicts the energy corresponding to the specified level of theory. [89]
HIP-NN Overview
All models in this study utilize the “Hierarchically Interacting Particle Neural Network with Tensor Sensitivity Information” (HIP-NN-TS) architecture, [38] an extension of the original HIP-NN architecture. [53] The original HIP-NN architecture is limited in terms of how information about atomic neighborhoods is collected by the model. The extended HIP-NN-TS architecture can incorporate either scalar, dipole, or quadrupole information about the feature distribution in the environment of an atom. HIP-NN-TS with tensor sensitivity of zero (i.e., scalar information) is equivalent to the original HIP-NN model. The accuracy (and computational cost) of HIP-NN-TS increases somewhat with increasing tensor sensitivity. To maximize accuracy, all HIP-NN-TS models reported in the main text utilize tensor sensitivity of $\ell=2$ (that is, using scalar, dipole, and quadrupole information). For comparison, Supplementary Material provides results for other HIP-NN-TS models with tensor sensitivity of zero (scalar information) and one (scalar and dipole information). For brevity, we refer to HIP-NN-TS simply as HIP-NN throughout the main text.
HIP-NN is a message-passing graph-based neural network that consists of different types of layers. At each layer, HIP-NN computes an atomic feature vector $z_{i}$ for each atom $i$. This atomic feature vector is an abstract numeric representation of the local chemical environment around atom $i$. $z_{i}$ is composed of $n_{\rm features}$ number of scalar components, $z_{i,a}$, for $a=1...n_{\rm features}$. The one exception is the input layer (see blue squares in Figure 4). The input feature vector encodes the atomic species (H, C, N, and O in this work) of each atom $i$ using one-hot encoding into a feature dimension indexed by $a$. Thus, the dimensionality of the initial feature vector is equal to the number of element types (four in this work).
Figure 4: (Left) Standard HIP-NN single-fidelity architecture with interaction blocks 1 to n. (Right) Interaction block containing one interaction layer followed by $n_{\rm layers}$ atomic layers.
A key feature of HIP-NN is the interaction block (see green rectangles in left panel of Figure 4). Interaction blocks contain one interaction layer followed by $n_{\rm layers}$ number of atomic layers (see right panel in Figure 4).
Interaction layers enable message passing between atoms within a predefined cut-off distance (6.5 Å in this work), mixing the feature vectors of neighboring atoms. Each interaction layer transforms the atomic feature vector, $z_{i}$, into an updated atomic feature vector, $z_{i}^{\prime}$, which is then passed to the next atomic layer. This transformation is given by
$z^{\prime}_{i,a}=f\left(I_{i,a}(z,r)+\sum_{b}W_{ab}z_{i,b}+B_{a}\right)$ (4)
where $W$ is a trainable weight matrix, $B$ is a trainable bias vector, and $f$ is a non-linear activation function. Both $W$ and $B$ operate locally for each atom $i$. The interaction term, $I$, provides the important information about the local atomic environment by incorporating contributions from neighboring atoms within the cut-off distance. The interaction term is defined as
$I_{i,a}(z,r)=\sum_{j,b}v_{ab}(r_{ij})z_{j,b}.$ (5)
Here, the interaction term aggregates information from the feature vectors $z_{j}$ of neighboring atoms $j$ (where $j\neq i$) within the cut-off distance. The function $v_{ab}$ depends on the interatomic distance $r_{ij}=|r_{i}-r_{j}|$, and is expanded as
$v_{ab}(r_{ij})=\sum_{\nu}V_{ab}^{\nu}s^{\nu}(r_{ij}),$ (6)
where $V^{\nu}$ is trainable weight matrix and $s^{\nu}(r_{ij})$ are a finite basis of trainable sensitivity functions.
Similar to interaction layers, atomic layers also apply a local transformation to the feature vector of each atom, $z_{i}$ to $z_{i}^{\prime}$. This transformation is given by:
$z^{\prime}_{i,a}=f\left(\sum_{n}w_{ab}z_{i,n}+b_{a}\right)$ (7)
where $w$ is a trainable weight matrix, $b$ is a trainable bias vector, and $f$ is a non-linear activation function.
In the spirit of many-body theory, HIP-NN adheres to the assumption that the total energy of a system can be separated into multiple hierarchical energy contributions. To this end, HIP-NN predicts a hierarchical energy for each atom $i$ (yellow boxes in Figure 4). This hierarchical energy is obtained from a linear regression of the atomic feature vector according to
$E^{n}_{i}=\sum_{a}w_{a}z_{i,a}+b$ (8)
where $E^{n}_{i}$ is the $n^{\rm th}$ hierarchical energy for atom $i$, $w$ is a trainable weight matrix and $b$ is a trainable bias vector. For $n=0$, $z_{i,a}$ corresponds to the one-hot encoding input feature vector. For $n>0$, $z_{i,a}$ corresponds to the transformed feature vector following the $n^{\rm th}$ interaction block. The hierarchical energies for a given atom $i$ are summed together to give the so-called atomic energy
$E_{i}=\sum_{n}E^{n}_{i}$ (9)
where, since all models in this work have only two interaction blocks, the sum is over $n=$ 0, 1, and 2. Finally, as is common practice for most MLIPs and classical FFs, these atomic energies are summed together over all atoms to obtain the total system energy
$E=\sum_{i}E_{i}$ (10)
To ensure conservation of energy, the forces acting on each atom are not predicted directly by HIP-NN. Instead, the forces are computed, using automatic differentiation, as the negative derivative of the total energy with respect to position for a given atom.